leslie buys 38 stamps for 9.88. If each stamp cost either 32 cents or 20 cents how much of each kind can she buy


answer: leslie bought nineteen 32 cents stamps and nineteen 20 cents stamps.
need a break down of how my text book got this answer

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(57347)

leslie buys 38 stamps for 9.88. If each stamp cost either 32 cents or 20 cents how much of each kind can she buy
-------------------------------------
let the 32 cents be "a" stamps
let the 20 cents be "b" stamps
-------------------------------------
Equations:
Quantity Eq. : a + b = 38
Value Eq.....:32a+20b= 988 cents
------------------------------------
Modify to set the system for elimination:
20a + 20b = 20*38
32a + 20b = 988
---------------------------------------
Subtract the 1st from the 2nd to solve for "a":
12a = 988 - 760
12a = 228
a = 12 (# of 32 cent stamps)
-------
Substitute to solve for "b":
12 + b = 38
b = 24 (# of 20 cent stamps)
================================
Cheers,
Stan H.

Answer by solver91311(16885)



Let x be the number of 32 cent stamps. Since there are 38 stamps total, the number of 20 cent stamps must be 38 - x.

Since 32 cent stamps cost 32 cents each, the amount she spent on 32 cent stamps must be 32x cents. Likewise, the amount spent on 20 cent stamps must be 20(38 - x). The total amount spent is $9.88, but since the amount for each kind of stamp is expressed in cents, we will change $9.88 to 988 cents. Now we can say:









John